6. Consider an economy without private banks (i.e. the Central Bank is the only bank).

Suppose that the following is true: L(i) = (.5-i)
$Y=P*Y=$10 trillion
Ms=$3.5 trillion

a. Graph money supply and money demand for this economy.
b. What is the equilibrium interest rate? Label this point on your graph from part (a).

c. Now, the central bank sells $1 trillion worth of bonds via open market operations. Find the new equilibrium interest rate. On your graph from part (a) illustrate the movement to this new equilibrium.

d. At the equilibrium from part (b), what is the price of a bond that promises to pay $100 at the end of a year?

e. At the equilibrium from part (c), what is the price of a bond that promises to pay $100 at the end of a year?

f. Why did the price of bonds fall?
Answer #1

6)a)

B)L(i)=0.5$Y-i

Ms=$3.5 trillion

Equilibrium

3.5=0.5*10-i

3.5=5-i

i=1.5%

C) New Ms=2.5 trillion

Equilibrium

2.5=0.5*10-i

i=5-2.5=2.5%

D)let that amount=x

X*1.015=100

X=100/1.015=98.522

E) let that amount=y

Y*1.025=100

Y=100/1.025=97.561

F)when interst rate increases ,then investors can get better rate of return from saving in bank account or from other assest. So because of this demand of bonds falls so it's price decline.